# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v1

@@ Abstract
Building generalist models has recently demonstrated remarkable capabilities in diverse scientific domains. Within the realm of molecular learning, several studies have explored unifying diverse tasks across diverse domains. However, negative conflicts and interference between molecules and knowledge from different domain may have a worse impact in threefold. First, conflicting molecular representations can lead to optimization difficulties for the models. Second, mixing and scaling up training data across diverse tasks is inherently challenging. Third, the computational cost of refined pretraining is prohibitively high. To address these limitations, this paper presents Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol builds on three key components to tackles conflicts: (1) a unified encoding mechanism for any task input; (2) an active-learning-driven data selection strategy that significantly reduces dataset size; (3) a novel design of the adaptive gradient stabilization module and anchor-and-reconcile MoE framework that ensures stable convergence. Experimentally, Omni-Mol achieves state-of-the-art performance across 15 molecular tasks, demonstrates the presence of scaling laws in the molecular domain, and is supported by extensive ablation studies and analyses validating the effectiveness of its design. The code and weights of the powerful AI-driven chemistry generalist are open-sourced at: https://anonymous.4open.science/r/Omni-Mol-8EDB.
@@ 1 Introduction
It drives us to develop Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol is the first generalist to effectively mitigate conflict collapse in three respects: (1) Omni-Mol proposes a unified encoding mechanism applicable to any task input, leveraging the most comprehensive instruction-following omni-molecular dataset to date, which comprises 1.8M samples across 15 tasks. (2) From the perspective of data, Omni-Mol employs an active learning-based dynamic data selection after recognizing that not all tasks are equally important. We significantly reduced the dataset size to 40% of its original volume while maintaining comparable performance with the full dataset. (3) From the perspective of architecture, our novel designed adaptive gradient stabilization successfully mitigate the rapid gradients growth caused by task conflicts (shown in Figure 2). Besides, our anchor-and-reconcile mixture-of-expert (MoE) architecture reduces interference by dynamically routing tasks to reconcile experts and capturing common knowledge through anchor experts.
Extensive experiments show that Omni-Mol achieves significant improvements across 15 tasks simultaneously, setting new state-of-the-art results among LLM-based models. Additionally, we observe that Omni-Mol scales effectively with increases in data volume, model size, and the number of tasks, indicating the model’s tremendous potential under larger computational budgets. Furthermore, by analyzing the representations of models trained with progressively more tasks, we discover that the representations become increasingly similar as the number of tasks grows. This provides robust evidence supporting our hypothesis of convergence space toward a universal molecular representation.
@@ 3.1 Overview
Omni-Mol is a multimodal LLM framework to handle $K$ diverse molecular tasks simultaneously. It comprises a language model, a graph encoder $f_{\mathcal{G}}$, and a projector $f_{p}$. The inputs include a text instruction $\mathbf{X}_{I}$, a SELFIES string $\mathbf{X}_{S}$, and the graph data $\mathbf{X}_{G}$ corresponding to the input molecules, where $\mathbf{X}_{G}$ is converted from $\mathbf{X}_{S}$ using RDKit. We model the response $\mathbf{Y}$ as the probability of the next token as: $P(\mathbf{Y}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G})=\prod_{i=1}^{L}P_{\theta}(\mathbf{Y}_{i}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G},\mathbf{Y}_{<i})$, where $\mathbf{H}_{G}=f_{p}(f_{\mathcal{G}}(\mathbf{X}_{G}))$, and $\theta$ is the parameter. The graph encoder encodes the molecule graph into its representation $\mathbf{h_{g}}\in\mathbb{R}^{n\times d_{1}}$, where $n$ is the length of the representation, the projector then projects its dimension to the LLM’s hidden size and obtain $\mathbf{H}_{G}\in\mathbb{R}^{n\times d_{2}}$.
@@ 3.3 Active Learning-based Data Selection
Given $K$ task collections $\mathcal{T}=\bigcup_{q=1}^{K}\mathcal{T}_{q}$, along with a total budget $\mathcal{B}$ and a maximum iteration count $J$. For the $j$-th iteration $\bigl{(}1\leq j\leq J\bigr{)}$, we select a data portion of size $\alpha_{j}=\mathcal{B}/L$. Assume the initial parameters of the model is $\theta_{0}$. For the $K$ subsets, we first initialize the distribution $\boldsymbol{\pi}^{0}=\bigl{(}\pi_{1}^{0},\,\pi_{2}^{0},\,\ldots,\,\pi_{K}^{0}\bigr{)},$ such that $\pi_{q}^{0}=1/K$, Then, we sample each task dataset $\mathcal{T}_{q}$ to a subset $\mathcal{T}_{q}^{(1)}$ according to the distribution and create the sampled training set $\mathcal{T}^{(1)}$. We then finetune $\theta_{0}$ on $\mathcal{T}^{(1)}$ for $M$ epochs, yielding the updated parameter $\theta_{1}$.
After obtaining $\theta_{j}$, we perform inference on the sampled dataset $\mathcal{T}^{(1)}$, we collect sample pair $(u_{p,+},u_{p,-})$ indexed by $p$, where $u_{p,+}$ is the ground truth and $u_{p,-}$ is the answer generated by the model. For each sample pair, we calculate the score of the sample by:
@@
We then select an additional batch of size $\alpha_{j+1}=\mathcal{B}/L$ denoted $\mathcal{T}^{(j+1)}$, from the unused portion of the dataset. We concatenate it with the previously chosen data for further training. This sample–train–evaluate–reweight routine continues until $J$ total iterations are reached.
@@ 3.4 Stable & MoE Expanded Framework
Let $s_{i}$ denote the output of the router logits for the $i$-th expert, assume $\mathcal{E}$ experts are chosen the output of the MoE layer can be written as:
$\mathbf{x_{t}^{\prime}}=\sum_{i=1}^{\mathcal{N}}\text{TopK}(s_{i},\mathcal{E})E_{\phi_{i}}(\mathbf{x_{t}})+E_{\phi_{\mathcal{N}+1}}(\mathbf{x_{t}})$ (8)
@@ 3.5 Optimization
For stage 2, we incorporate an additional auxiliary load balancing loss for the MoE layers, assume an input tensor $x\in\mathbb{R}^{B\times T\times d}$, and Top-$K$ experts out of $\mathcal{N}$ is selected, the load balancing loss is: $\mathcal{L}_{\text{aux}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{\mathcal{N}}C_{ij}\cdot\bar{s}_{ij}$, where $C_{ij}=\frac{\mathcal{N}}{T\mathcal{E}}\sum_{t=1}^{T\mathcal{E}}\mathbf{1}\{\text{t'th token selects expert j}\},s_{ij}=\frac{1}{T}\sum_{t=1}^{T}s_{i,j,t}$ and $\mathbf{1}\{\cdot\}$ is an indicator function. This load balancing loss used in Liu et al. (2024a) additionally considers the sequence-level information.
@@ 3.6 Theoretical Analysis
In our experimental validation, we construct a series of mixed datasets comprising 1, 2, 4, and 8 tasks to form the solution space sequence $S_{\mathcal{F}}$. on each of these datasets and subsequently extract representations for a specific task. Finally, we calculate the similarity between representations $R_{i}$ and $R_{j}$ for $i,j=0,\dots,3$. See Appendix C.5 for the calculation of similarity.
@@ 4.1 Setup and Baselines
  • Regression Tasks: LogP Prediction ($\sim$10k), SCF Prediction ($\sim$623k), TPSA Prediction ($\sim$11k), Complecity Prediction ($\sim$11k), Molecular Weight Prediction ($\sim$11k), Yield Regression ($\sim$9k) from, HOMO LUMO Property Prediction ($\sim$36k).
@@ 4.2 Main Results
Figure 4: Scaling trend of Omni-Mol. (Left) The scaling trend respect to dataset proportion, metrics are averaged across tasks, (Right) The scaling trend respect to model size, the metrics of HOMO LUMO task are normalized. we observed a clear log scaling behavior.
@@
HOMO LUMO Prediction Task
@@
Table 3: Main results of HOMO LUMO prediction task, 1: In-Context Learning results from Fang et al. (2024), 2: 5-shot In-Context Learning results from Cao et al. (2025).
@@ 4.3 Is Omni-Mol a Scalable Framework?
One critical property of Large Language Models (LLMs) is their scaling behavior in relation to both model and data size. In this study, we demonstrate that Omni-Mol is a scalable framework by conducting three distinct types of scaling experiments: (1) We select three different sizes of LLMs from the LLaMA 3 series, 1B, 3B, and 8B, for language backbone scaling. (2) We evaluate the impact of dataset size by down-sampling the original dataset to 20%, 40%, 60%, and 100% of its full size. (3) To examine task scaling, we train Omni-Mol on different numbers of tasks, specifically, 1, 2, 4, 8, and 15 tasks, and observe the performance of individual tasks within these multi-task settings.
@@ 5 Conclusion
Limitation We identified two limitations: (1) Due to the limited computational resources, we are unable to further scale up the model with higher computational resources, which prevents us from exploring the limits of model’s performance. (2) Omni-Mol’s tasks are still primarily focused on molecular understanding, and it is not yet capable of designing new molecules. Future work should explore unifying understanding and generation.
@@ B.1 Comprehensive Datasets Construction
(1) USPTO (USPTO, 2020). The USPTO (United States Patent and Trademark Office) dataset is a widely used large-scale chemical reaction dataset extracted and processed from US patent texts. It encompasses a diverse range of organic reaction types, including esterification, amidation, halogenation, Suzuki coupling, Buchwald–Hartwig coupling, addition reactions, condensation reactions, and redox reactions. Following Fang et al. (2024), for the Forward Reaction Prediction task, we extract data from USPTO, and split the dataset into 124,384 training instances and 1,000 test instances. Partially following Cao et al. (2024), for the Catalyst Prediction and Solvent Prediction tasks, we similarly extract data from USPTO, splitting the training/validation/test sets into 10,079/1,015 and 67,099/7,793, respectively.
USPTO-500-MT (Lu & Zhang, 2022) is a high-quality multi-task reaction prediction dataset, derived from USPTO through manual processing (including data filtering, deduplication, etc.). This subset retains the 500 most common reaction types. Following Fang et al. (2024), for the Reagent Prediction task, we split the dataset into 124,384 training instances and 1,000 test instances. For the Retrosynthesis task, the dataset is divided into 128,684 training instances and 1,000 test instances.
@@ Appendix D Further Details on Training
Separate instruction tuning. We follow the training recipe outlined in Cao et al. (2025). However, we observe significant overfitting when training the model on the molcap task for 20–50 epochs, as suggested in Cao et al. (2025). To address this issue, we manually allocate $10\%$ of the training set for validation and re-evaluated all tasks, we find that the recipes for forward prediction, reagent prediction, retrosynthesis, HOMO LUMO prediction from the original paper matches our results, however, we identify an updated training strategy tailored to the molcap task. The revised training recipe is summarized in Table 4.

### v2

@@ Abstract
Building generalist models has recently demonstrated remarkable capabilities in diverse scientific domains. Within the realm of molecular learning, several studies have explored unifying diverse tasks across diverse domains. However, negative conflicts and interference between molecules and knowledge from different domain may have a worse impact in threefold. First, conflicting molecular representations can lead to optimization difficulties for the models. Second, mixing and scaling up training data across diverse tasks is inherently challenging. Third, the computational cost of refined pretraining is prohibitively high. To address these limitations, this paper presents Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol builds on three key components to tackles conflicts: (1) a unified encoding mechanism for any task input; (2) an active-learning-driven data selection strategy that significantly reduces dataset size; (3) a novel design of the adaptive gradient stabilization module and anchor-and-reconcile MoE framework that ensures stable convergence. Experimentally, Omni-Mol achieves state-of-the-art performance across most of the 15 molecular tasks, demonstrates the presence of scaling laws in the molecular domain, and is supported by extensive ablation studies and analyses validating the effectiveness of its design. The dataset, code and weights of the powerful AI-driven chemistry generalist are open-sourced.
@@
[Uncaptioned image]Code:Omni-Mol-Code      [Uncaptioned image]Data:Omni-Mol-Data
@@ 1 Introduction
This question drives us to develop Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol is the first generalist to effectively mitigate conflict collapse in three respects: (1) Omni-Mol proposes a unified encoding mechanism applicable to any task input, leveraging the most comprehensive instruction-following omni-molecular dataset to date, which comprises 1.8M samples across 15 tasks. (2) From the perspective of data, Omni-Mol employs an active learning-based dynamic data selection after recognizing that not all tasks are equally important. We significantly reduced the dataset size to 40% of its original volume while maintaining comparable performance with the full dataset. (3) From the perspective of architecture, our novel designed adaptive gradient stabilization successfully mitigate the rapid gradients growth caused by task conflicts (shown in Figure 2). Besides, our anchor-and-reconcile mixture-of-expert (MoE) architecture reduces interference by dynamically routing tasks to reconcile experts and capturing common knowledge through anchor experts.
Extensive experiments show that Omni-Mol achieves significant improvements across most of the tasks, setting new state-of-the-art results among LLM-based models. Additionally, we observe that Omni-Mol scales effectively with increases in data volume, model size, and the number of tasks, indicating the model’s tremendous potential under larger computational budgets. Furthermore, by analyzing the representations of models trained with progressively more tasks, we discover that the representations become increasingly similar as the number of tasks grows. This provides robust evidence supporting our hypothesis of convergent space toward a universal molecular representation.
@@ 3.1 Overview
Omni-Mol is a multimodal LLM framework to handle $K$ diverse molecular tasks simultaneously. It comprises a language model, a graph encoder $f_{\mathcal{G}}$, and a projector $f_{p}$. The inputs include a text instruction $\mathbf{X}_{I}$, a SELFIES string $\mathbf{X}_{S}$, and the graph data $\mathbf{X}_{G}$ corresponding to the input molecules, where $\mathbf{X}_{G}$ is converted from $\mathbf{X}_{S}$ using RDKit. We model the response $\mathbf{Y}$ as the probability of the next token as:
$P(\mathbf{Y}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G})$ (1)
$=\prod_{i}P_{\theta}(\mathbf{Y}_{i}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G},\mathbf{Y}_{<i})$
where $\mathbf{H}_{G}=f_{p}(f_{\mathcal{G}}(\mathbf{X}_{G}))$, and $\theta$ is the parameter of the LLM. The graph encoder encodes the molecule graph into its representation $\mathbf{h_{g}}\in\mathbb{R}^{n\times d_{1}}$, where $n$ is the length of the representation, the projector then projects its dimension to the LLM’s hidden size and obtain $\mathbf{H}_{G}\in\mathbb{R}^{n\times d_{2}}$.
@@ 3.3 Active Learning-based Data Selection
Given $K$ task collections $\mathcal{T}=\bigcup_{q=1}^{K}\mathcal{T}_{q}$, along with a total budget $\mathcal{B}$ and a maximum iteration count $J$. For the $j$-th iteration $\bigl{(}1\leq j\leq J\bigr{)}$, we select a data portion of size $\alpha_{j}=\mathcal{B}/J$. Assume the initial parameters of the model is $\theta_{0}$. For the $K$ subsets, we first initialize the distribution $\boldsymbol{\pi}^{0}=\bigl{(}\pi_{1}^{0},\,\pi_{2}^{0},\,\ldots,\,\pi_{K}^{0}\bigr{)},$ such that $\pi_{q}^{0}=1/K$, Then, we sample each task dataset $\mathcal{T}_{q}$ to a subset $\mathcal{T}_{q}^{(1)}$ according to the distribution and create the sampled training set $\mathcal{T}^{(1)}$. We then finetune $\theta_{0}$ on $\mathcal{T}^{(1)}$ for $M$ epochs, yielding the updated parameter $\theta_{1}$.
After obtaining $\theta_{1}$, we perform inference on the sampled dataset $\mathcal{T}^{(1)}$, we collect sample pair $(u_{p,+},u_{p,-})$ indexed by $p$, where $u_{p,+}$ is the ground truth and $u_{p,-}$ is the answer generated by the model. For each sample pair, we calculate the score of the sample by:
@@
We then select an additional batch of size $\alpha_{j+1}=\mathcal{B}/J$ denoted $\mathcal{T}^{(j+1)}$, from the unused portion of the dataset. We concatenate it with the previously chosen data for further training. This sample–train–evaluate–reweight routine continues until $J$ total iterations are reached.
@@ 3.4 Stable & MoE Expanded Framework
Let $s$ denote the output of the router logits for an input token, assume $\mathcal{E}$ experts are chosen, the output of the MoE layer can be written as:
$\mathbf{x_{t}^{\prime}}=\sum_{i=1}^{\mathcal{N}}\text{TopK}(s,\mathcal{E})E_{\phi_{i}}(\mathbf{x_{t}})+E_{\phi_{\mathcal{N}+1}}(\mathbf{x_{t}})$ (9)
@@ 3.5 Optimization
For stage 2, we incorporate an additional auxiliary load balancing loss for the MoE layers, assume an input tensor $x\in\mathbb{R}^{B\times T\times d}$, and $\mathcal{E}$ experts out of $\mathcal{N}$ is selected, the load balancing loss is: $\mathcal{L}_{\text{aux}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{\mathcal{N}}C_{ij}\cdot\bar{s}_{ij}$, where $C_{ij}=\frac{\mathcal{N}}{T\mathcal{E}}\sum_{t=1}^{T\mathcal{E}}\mathbf{1}\{\text{t'th token selects expert j}\},\bar{s}_{ij}=\frac{1}{T}\sum_{t=1}^{T}s_{i,j,t}$ and $\mathbf{1}\{\cdot\}$ is an indicator function. This load balancing loss used in Liu et al. (2024a) additionally considers the sequence-level information.
@@ 3.6 Theoretical Analysis
In our experimental validation, we construct a series of mixed datasets comprising 1, 2, 4, and 8 tasks to form the solution space sequence $S_{\mathcal{F}}$. We train the model on each of these datasets and subsequently extract representations for a specific task. Finally, we calculate the similarity between representations $R_{i}$ and $R_{j}$ for $i,j=0,\dots,3$. See Appendix C.5 for the calculation of similarity.
@@ 4.1 Setup and Baselines
  • Regression Tasks: LogP Prediction ($\sim$10k), SCF Prediction ($\sim$623k), TPSA Prediction ($\sim$11k), Complecity Prediction ($\sim$11k), Molecular Weight Prediction ($\sim$11k), Yield Regression ($\sim$9k) from, Quantum Mechanics Property Prediction ($\sim$36k).
@@ 4.2 Main Results
Figure 4: Scaling trend of Omni-Mol. (Left) The scaling trend respect to dataset proportion, metrics are averaged across tasks, (Right) The scaling trend respect to model size, the metrics of Quantum Mechanics Property Prediction task are normalized. we observed a clear log scaling behavior.
@@
Quantum Mechanics Property Prediction Task
@@
Table 3: Main results of Quantum Mechanics Property Prediction task, 1: In-Context Learning results from Fang et al. (2024), 2: 5-shot In-Context Learning results from Cao et al. (2025).
@@ 4.3 Is Omni-Mol a Scalable Framework?
One critical property of LLMs is their scaling behavior in relation to both model and data size. In this study, we demonstrate that Omni-Mol is a scalable framework by conducting three distinct types of scaling experiments: (1) We select three different sizes of LLMs from the LLaMA 3 series, 1B, 3B, and 8B, for language backbone scaling. (2) We evaluate the impact of dataset size by down-sampling the original dataset to 20%, 40%, 60%, and 100% of its full size. (3) To examine task scaling, we train Omni-Mol on different numbers of tasks, specifically, 1, 2, 4, 8, and 15 tasks, and observe the performance of individual tasks within these multi-task settings.
@@ 5 Conclusion
Limitation. We identified two limitations: (1) Due to the limited computational resources, we are unable to further scale up the model with higher computational resources, which prevents us from exploring the limits of model’s performance. (2) Omni-Mol’s tasks are still primarily focused on molecular understanding, and it is not yet capable of designing new molecules. Future work should explore unifying understanding and generation.
@@ B.1 Comprehensive Datasets Construction
(1) USPTO (USPTO, 2020). The USPTO (United States Patent and Trademark Office) dataset is a widely used large-scale chemical reaction dataset extracted and processed from US patent texts. It encompasses a diverse range of organic reaction types, including esterification, amidation, halogenation, Suzuki coupling, Buchwald–Hartwig coupling, addition reactions, condensation reactions, and redox reactions. Following Fang et al. (2024), for the Forward Reaction Prediction task, we extract data from USPTO, and split the dataset into 124,384 training instances and 1,000 test instances. Partially following Cao et al. (2024), for the Catalyst Prediction and Solvent Prediction tasks, we similarly extract data from USPTO, splitting the training/test sets into 10,079/1,015 and 67,099/7,793, respectively.
USPTO_500_MT (Lu & Zhang, 2022) is a high-quality multi-task reaction prediction dataset, derived from USPTO through manual processing (including data filtering, deduplication, etc.). This subset retains the 500 most common reaction types. Following Fang et al. (2024), for the Reagent Prediction task, we split the dataset into 124,384 training instances and 1,000 test instances.
USPTO_500K (Lu & Zhang, 2022), a subset of organic chemical reaction data extracted from USPTO, is widely used in chemoinformatics for the single-step retrosynthesis task. Following Fang et al. (2024), for the Retrosynthesis task, the dataset is divided into 128,684 training instances and 1,000 test instances.
@@ Appendix D Further Details on Training
Separate instruction tuning. We follow the training recipe outlined in Cao et al. (2025). However, we observe significant overfitting when training the model on the molcap task for 20–50 epochs, as suggested in Cao et al. (2025). To address this issue, we manually allocate $10\%$ of the training set for validation and re-evaluated all tasks, we find that the recipes for forward prediction, reagent prediction, retrosynthesis, Quantum Mechanics Property Prediction from the original paper matches our results, however, we identify an updated training strategy tailored to the molcap task. The revised training recipe is summarized in Table 4.